OpenStudy (anonymous):

the sum of two numbers is 23. the larger of the two numbersis 1 less than twice the smaller. what is the difference between the larger and the smaller number?

OpenStudy (anonymous):

let the larger number be x and smaller be y Therefore, x = 2y - 1 and their sum = 23, x+y = 23 => 2y-1+y = 23 => 3y = 24 => y = 8 => x = 23-y = 23-8 = 15 Therefore, difference = difference between 8 and 15, which is 7

OpenStudy (anonymous):

thank you :)